# Nowcast3D: Reliable precipitation nowcasting via gray-box learning

### v3

@@ Introduction
The inferred velocity field represents the motion of hydrometeors and is therefore most reliable in regions with substantial radar reflectivity. Its accuracy is inherently limited in echo-free areas and where low-level radar coverage is sparse. Within the precipitating volume, however, the model yields a dynamically consistent, high-resolution wind field that complements the spatially sparse measurements from ground-based stations. The calculation and evaluation of this wind field are summarized in Extended Data Figure 5, and the velocity field derived from the stream-function and potential-function decomposition is detailed in Supplementary Note 6.2.
@@
Figure 3: Generalization of Nowcast3D to high resolution, fine scale forecasting. a, Radar reflectivity over Maoming, Guangdong Province, at the forecast initialization time. The analysis was conducted at 0.01° resolution. b, Quantitative forecast skill comparison using the CSI at 30 dBZ and 40 dBZ reflectivity thresholds over the 3 hour forecast horizon. c, Analysis of spatial scale fidelity using PSD. Forecast spectra are compared to the observed spectrum (black dashed line). d, Perceptual similarity assessment using the LPIPS metric. Lower values indicate better performance. e, Probabilistic forecast skill assessed using the CRPS, shown as both average-pooled and max-pooled values. Lower scores indicate better-calibrated ensemble forecasts. f, Vertical profiles of forecast error and skill. MAE and LPIPS are plotted as a function of height, overlaid with the mean observed reflectivity and data coverage profiles. g, Visual comparison of forecast reflectivity fields from Nowcast3D and baseline models against the observed radar sequence.
@@
To assess the model’s ability to generalize to finer spatial scales, which is a key requirement for urban and operational forecasting, we evaluate its performance on an extreme precipitation event in Maoming on 16 August 2025. For this case, all models are run at a 0.01° spatial resolution (Figure 3a). This higher resolution resolves faster and more intricate atmospheric processes and therefore poses a demanding test of forecast skill.
@@
The storm evolution was dominated by two convective cells moving eastward along the coast. The western cell, initially more compact, broadened markedly and intensified as it progressed, whereas the eastern cell translated eastward and gradually moved out of the forecast domain. The key forecasting challenge was therefore to reproduce the changing storm morphology under rapid advection, in particular the rapid area growth of the western cell and the simultaneous displacement of the eastern cell, which together produced large changes in the reflectivity field.
@@
Comparing the predicted spatiotemporal evolution at this high resolution highlights the difficulty of forecasting such rapidly changing radar fields (Figure 3g). The deep-learning baselines, NowcastNet and SimVP, reproduce the overall eastward motion but fail to capture the expansion of the western cell, leading to fragmented and underdeveloped echoes. The parameter-free pySTEPS model underestimates the translation speed of the system and does not reproduce its area growth. In contrast, Nowcast3D more faithfully captures both the rapid eastward displacement and the structural expansion of the convective system. Quantitative results support these visual assessments. Nowcast3D outperforms all baseline models by a wide margin in the neighbourhood CSI, indicating more accurate localization of precipitation (Figure 3b). In addition, PSD and LPIPS evaluations show that its forecasts better reproduce the observed scale-dependent variability and exhibit higher perceptual similarity to the radar observations (Figure 3c,d). Because Nowcast3D produces ensembles, we also assess probabilistic skill. The Continuous Ranked Probability Score (CRPS) [matheson1976scoring, hersbach2000decomposition] shows that Nowcast3D yields better-calibrated forecasts than the ensemble-based NowcastNet, indicating a more reliable quantification of uncertainty (Figure 3e).
@@
Taking advantage of its native three-dimensional design, we further examine the vertical distribution of forecast error (Figure 3f). A layer-wise analysis of absolute error shows smaller values near the surface and at upper levels, and larger values in mid-tropospheric layers. This pattern closely follows the vertical profile of mean reflectivity: echoes are strongest and most structured at mid-levels, leading to larger absolute differences, whereas near-surface and upper levels exhibit weaker and more intermittent echoes, which naturally yield smaller absolute errors. The vertical error profile therefore reflects both the intrinsic storm structure and the sampling characteristics of the radar, rather than a simple ranking of forecast skill with height. Layer-by-layer comparisons of the three-dimensional forecasts are provided in Extended Data Figure 4.
@@
Figure 4: Aggregate quantitative evaluation of forecast skill. a, Information of participating meteorologists, including gender distribution and regional distribution of their affiliated institutions. b, Evaluation of meteorologists’ model preferences across different regions of China. Compared with other competing methods, meteorologists showed a stronger preference for Nowcast3D. c, Forecast skill metrics evaluated on the 0.04° resolution test set. The top panels show the CSI as a function of lead time at 35 dBZ and 45 dBZ thresholds. The bottom-left panel shows the PSD of forecasts compared with the ground truth (black dashed line), and the bottom-right panel shows the mean LPIPS, where lower values indicate better perceptual quality. d, The same set of metrics as in c, evaluated on the 0.01° resolution test set to assess generalization at finer spatial scales.
@@
To assess the usefulness of different models for forecasting extreme precipitation, we conducted an expert evaluation following the protocols of DGMR [ravuri2021skilful] and NowcastNet [zhang2023skilful]. A nationwide invitation was issued to senior forecasters across China. Forecast animations from all competing models were displayed side by side, anonymized and presented in random order to minimize bias.
In total, 160 meteorologists from the China Meteorological Administration and 30 provincial or municipal observatories participated. The demographic and regional distribution of participants is shown in Figure 4a. The experts cover almost all regions of China, except Tibet, Hong Kong, Macao and Taiwan. Each meteorologist evaluated 23 test cases, randomly sampled from precipitation events that occurred over China. Following the NowcastNet protocol [zhang2023skilful], we considered two complementary evaluation modes. In the post-hoc mode, forecasters ranked the model forecasts with access to the verifying radar observations. In the prior mode, they ranked the forecasts using only the preceding radar sequences, without access to future outcomes, to mimic real-time operational decision-making.
Aggregated preferences are summarized in Figure 4b. In the post-hoc evaluation, Nowcast3D was ranked as the best forecast in 57% of cases; in the prior evaluation, it remained the top choice in 51% of cases. Overall, Nowcast3D received the highest preference among participating meteorologists. These preferences are consistent with the model’s ability to represent the three-dimensional evolution of reflectivity, enabled by its explicit treatment of physical fields and embedded dynamical constraints. By contrast, models based on two-dimensional formulations showed limitations in reproducing the full spatiotemporal structure of precipitation systems. A detailed description of the expert evaluation configuration is provided in Supplementary Note 4.
@@
To quantify forecast skill, we use three complementary metrics: the neighbourhood CSI to assess location-dependent event detection, the PSD to evaluate scale-dependent spatial structure, and LPIPS to measure perceptual and structural similarity. All metrics are computed over the full test corpus at two spatial scales: a provincial domain at a standard 0.04° resolution and a more demanding urban domain at 0.01° resolution. This dual-scale design allows us to examine both baseline performance and robustness to finer spatial detail.
At 0.04° resolution, Nowcast3D consistently outperforms all baseline models across all metrics (Figure 4c). For both moderate (35 dBZ) and intense (45 dBZ) reflectivity thresholds, it attains the highest neighbourhood CSI throughout the three-hour forecast horizon, indicating more accurate localization of heavy precipitation. PSD analysis shows that its forecasts most closely match the observed power spectrum, reflecting a more realistic representation of precipitation structures across spatial scales. This is corroborated by the LPIPS scores, for which Nowcast3D achieves the lowest values, indicating the highest perceptual and structural fidelity.
This advantage persists at 0.01° resolution, which provides a stricter test of performance at city scales (Figure 4d). Across all lead times and metrics, Nowcast3D maintains a clear margin over the competing methods, demonstrating that its architecture can be effectively adapted to finer spatial grids. Taken together, these results show that explicitly modeling three-dimensional dynamics yields forecasts that are more accurate, structurally realistic and perceptually coherent than those produced by existing approaches, and that this benefit is robust across spatial scales.
Additional CSI curves over a wider range of intensity thresholds are provided in Supplementary Note 6.1, Figure 5, and additional precipitation case studies together with comparisons of forecast results are presented in Supplementary Note 6.3.
@@ Discussion
Precipitation nowcasting remains a central challenge in atmospheric science, governed by intrinsically stochastic and multiscale processes. Recent advances in data-driven methods have improved short-range guidance, yet predictive skill for high-impact weather still degrades rapidly with lead time. This limitation is particularly severe for events in which complex vertical structure and microphysical evolution play a dominant role. Methods that rely on two-dimensional data neglect the vertical dimension, where key processes such as vertical motion and wind shear unfold, and therefore struggle to achieve sufficient physical fidelity and operational value in hazardous weather regimes.
@@
Our work addresses these limitations by combining physics-informed dynamics with deep generative modeling on native three-dimensional radar volumes. The central idea is a model that learns to infer and extrapolate key physical drivers directly from sequences of volumetric reflectivity: a three-dimensional wind field that governs advective transport, a local field that represents stochastic dispersion, and a residual tendency associated with unresolved microphysics. The forecast is advanced according to these learned operators, enforcing dynamical consistency and enabling physically plausible predictions at extended lead times. This physics-driven core is coupled to a conditional generative model that approximates the full distribution of future states, capturing uncertainty in both storm structure and localized intensity. Together, these components yield forecasts that are dynamically consistent and probabilistically reliable, providing the uncertainty information required for risk-aware decision-making.
@@
The performance of this framework nonetheless depends on several external factors. First, its accuracy is constrained by the quality, coverage and scan strategy of the input radar network, and beam attenuation, ground clutter and anomalous propagation remain important sources of error. Second, inferring three-dimensional flow from reflectivity alone is underconstrained, particularly in regimes with complex microphysics where phase changes can alter reflectivity without a corresponding change in the velocity field. These limitations point to clear directions for future work. Integrating additional observation types, such as Doppler radial velocities, geostationary satellite imagery and lightning measurements, could better constrain the inferred wind fields and improve the realism of hydrometeor evolution. A closer coupling with numerical weather prediction models, for example through data assimilation of the learned fields or by using NWP output as a physical prior, could further extend forecast skill beyond the traditional nowcasting range.
@@
In conclusion, by unifying physics-informed dynamics with deep generative modeling on three-dimensional radar data, this study outlines a pathway toward a new generation of nowcasting systems. Such models move beyond simple pattern extrapolation to learn representations of the underlying physical processes, producing forecasts that are more accurate, reliable and physically consistent. Further enriching this framework with multimodal observations and tighter NWP coupling will be essential for closing the remaining gap toward robust, actionable forecasts of high-impact weather that are urgently needed in practice.
@@ Methods
Nowcast3D is a hybrid forecasting system that combines a physics-informed predictor with a diffusion-based generative model to produce probabilistic three-dimensional radar reflectivity forecasts. As summarised in Extended Data Figure 1, the workflow has three stages. First, raw volumetric radar observations are passed through a data-processing module that derives a three-dimensional spatial mask of observed and unobserved regions, applies mask-guided imputation only in the unobserved regions, and normalises reflectivity to a fixed dynamic range. Second, a physics-informed network infers underlying physical fields from the recent radar history and advances them forward in time through an explicit advection–diffusion–source equation, yielding a deterministic forecast. Third, a conditional diffusion model refines this physics-consistent forecast into an ensemble of stochastic realisations that quantify forecast uncertainty.
In this framework, the preprocessing stage defines the input reflectivity fields $R$ used throughout the paper. The physics-informed module implements a mapping $\boldsymbol{\Theta}$ from radar history to time-varying physical fields $X_{1:T}$ and a numerical solver for the resulting evolution operator $\mathcal{F}$. The generative module implements a conditional diffusion model $\boldsymbol{\Phi}$ that samples future trajectories conditioned on both past observations and the deterministic forecast. The following subsections describe the radar data and preprocessing, the hybrid physics–generative formulation, the physics-informed forecasting module, the probabilistic refinement module, and the training procedure.
@@
Our forecasts are driven solely by three-dimensional radar reflectivity from the national radar composite operated by the China Meteorological Administration. This product provides volumetric reflectivity on 24 constant-altitude levels between 0.5 and 16 km, with native horizontal resolutions down to 0.01°. These volumes resolve the three-dimensional structure and evolution of precipitating systems over China and form the common input to all models evaluated in this study. Details of the dataset, including the storage format, are provided in Supplementary Note 1.1.
We consider three study domains that together span diverse convective regimes and spatial resolutions. The primary training domain is South China (110.00°–120.24°E, 23.96°–34.20°N) in 2024, where we use all eligible radar volumes. Cross-regional generalisation at the native 0.04° resolution is assessed on a subset of precipitating scenes over North China (110.00°–120.24°E, 33.96°–44.20°N) drawn from 2025; this subset is spatially and temporally disjoint from the South China training domain. High-resolution coastal hazards are evaluated in a 0.01° experiment over the Maoming region in Guangdong Province (109.64°–112.20°E, 20.39°–22.95°N), where models are first trained on South China and then fine-tuned on 0.01° Maoming sequences from 2024; an independent subset of Maoming precipitation events from 2025 is reserved for testing. The radar composite coverage and the three study domains are shown in Extended Data Figure 2a, and Extended Data Figure 2b illustrates how terrain blocking and the shorter range of low-elevation scans lead to a pattern of sparse, near-surface circular observations.
Before entering the forecasting model, raw radar volumes pass through a preprocessing pipeline (Data Process in Extended Data Figure 1b). We first construct a mask that flags regions with missing measurements, including areas affected by terrain-induced beam blockage and range limitations. A non-learning, mask-guided imputation scheme is then applied to fill only the masked regions, using upper-level reflectivity as a predictor while leaving all observed echoes unchanged. Finally, reflectivity values are normalised to a fixed dynamic range. Extended Data Figure 2c,d compare raw and imputation reflectivity fields at 500 m and 1,000 m over the South China and Maoming domains, respectively. (see Supplementary Notes 1.2–1.4 for further details on masking, imputation and normalisation).
@@
Let $R\in\mathbb{R}^{T\times D\times H\times W}$ denote a sequence of preprocessed three-dimensional radar reflectivity fields, where $T$ is the number of lead times and $D\times H\times W$ are the vertical and horizontal grid dimensions. The past observations are written as $R_{-T_{0}:0}\in\mathbb{R}^{T_{0}\times D\times H\times W}$. Given $R_{-T_{0}:0}$, our goal is to forecast the future sequence $R_{1:T}$ using a two-stage framework that first produces a physics-based deterministic forecast and then refines it with a probabilistic ensemble model.
In the first, physics-driven stage, a neural network $\boldsymbol{\Theta}$ maps the historical radar sequence to a set of time-varying physical fields, $X_{1:T}=\boldsymbol{\Theta}(R_{-T_{0}:0})$, where $X_{1:T}\in\mathbb{R}^{T\times C\times D\times H\times W}$ and $C$ is the number of inferred quantities that govern the evolution of reflectivity. These fields act as coefficients in a time-dependent evolution equation, $\partial_{t}R=\mathcal{F}(R,X)$, where $\mathcal{F}$ represents the physical dynamics controlled by $X$. Starting from the initial condition $R(0)=R_{0}$ (the last observed volume), integrating this equation forward in time yields a deterministic forecast $R^{\mathrm{det}}_{1:T}$.
In the second, probabilistic stage, a conditional diffusion model generates an ensemble of forecasts around this physics-consistent trajectory. The generative network $\boldsymbol{\Phi}$ is conditioned on the context $c=\{R_{-T_{0}:0},R^{\mathrm{det}}_{1:T}\}$, which combines the past observations with the deterministic prediction. Ensemble members are obtained by sampling the learned conditional distribution with different noise realisations $\epsilon^{(k)}$, $k=1,\dots,K$, thereby representing stochastic variability and forecast uncertainty.
These steps can be written compactly as
@@
Nowcast3D implements this formulation with a shared backbone that realises $\boldsymbol{\Theta}$ and a conditional diffusion network that realises $\boldsymbol{\Phi}$, as described next.
@@
The deterministic forecast $R^{\mathrm{det}}_{1:T}$ is obtained by numerically integrating a partial differential equation. We specify the dynamics $\mathcal{F}(R,X)$ as an advection–diffusion–source equation:
@@
Here the reflectivity field $R$ evolves under three processes: advective transport by a three-dimensional velocity field $\boldsymbol{v}$; local anisotropic diffusion governed by a positive-definite diffusivity tensor $\boldsymbol{\kappa}$; and a residual source term $s$ that summarises unresolved processes, including microphysical growth and decay. The time-varying fields $(\boldsymbol{v},\boldsymbol{\kappa},s)$ over the forecast horizon constitute the physical fields $X_{1:T}$ and are inferred from the input radar sequence $R_{-T_{0}:0}$ by the physics module $\boldsymbol{\Theta}$.
Neural parameterisation.
The time-varying physical fields $X_{1:T}=(\boldsymbol{v}_{1:T},\boldsymbol{\kappa}_{1:T},s_{1:T})$ required in Equation (2) are produced by a shared encoder–decoder backbone followed by task-specific heads. We denote the backbone (encoder, U-Net branch, Transformer branch and fusion) by $\mathrm{Backbone}(\cdot)$. Its output is passed to three dedicated decoders that specialize in motion, diffusion and source components:

### v4

@@ Introduction
The inferred velocity field represents the motion of hydrometeors and is therefore most reliable in regions with substantial radar reflectivity. Its accuracy is inherently limited in echo-free areas and where low-level radar coverage is sparse. Within the precipitating volume, however, the model produces a dynamically consistent, high-resolution wind field that complements the spatially sparse measurements by ground-based stations. The calculation and evaluation of this wind field are summarized in Extended Data Figure 4, and the velocity field derived from the stream-function and potential-function decomposition is detailed in Supplementary Note 6.2. Another precipitation case study at a different time, together with forecast comparisons, is presented in Supplementary Note 6.3 (see Supplementary Figure 9 and Supplementary Video 2).
@@
Figure 3: Generalization of Nowcast3D to high-resolution, fine-scale forecasting ($T$ = 13:54:$0$ pm, June 17, 2025). a, Radar reflectivity over Maoming, Guangdong Province, China at the forecast initialization time. The analysis was conducted at 0.01° resolution. b, Quantitative comparison of forecast skills in context of CSI at 30 dBZ and 40 dBZ reflectivity thresholds over a 3-hour forecast horizon. c, Analysis of spatial scale fidelity based on PSD. The forecast spectra are compared to the observed spectrum (black dashed line). d, Perceptual similarity assessment based on the LPIPS metric. Lower values indicate better performance. e, Probabilistic forecast skill evaluated using the CRPS metric, shown as both average-pooled and max-pooled values. Lower scores indicate better-calibrated ensemble forecasts. f, Vertical profiles of forecast error and skill. CSI and LPIPS are plotted as a function of height, overlaid with profiles of mean observed reflectivity and data coverage. g, Forecast reflectivity fields by Nowcast3D and baseline models, compared with the observed radar sequence. The frame-by-frame comparison at 6-min intervals is provided in Supplementary Video 3.
@@
To assess our model’s ability to generalize to finer spatial scales (a key requirement for urban and operational forecasting), we evaluate its performance on forecasting an extreme precipitation event in Maoming, Guangdong Province, China on August 16, 2025. In this case, all models are operated at a 0.01° spatial resolution (Figure 3a). This higher resolution resolves faster and more intricate atmospheric processes, thereby posing a demanding test of forecast skill.
@@
The storm evolution was dominated by two convective cells moving eastward along the coast. The western cell, initially more compact, broadened and intensified significantly as it progressed, while the eastern cell translated eastward and gradually moved out of the forecast domain. The key forecasting challenge was thus to reproduce the evolving storm morphology under rapid advection, in particular, the rapid area growth of the western cell and the simultaneous displacement of the eastern cell, which together drove large changes in the reflectivity field.
@@
Comparing the predicted spatiotemporal evolution at this high resolution reveals the challenge of forecasting such rapidly evolving radar fields (Figure 3g). The deep-learning baselines, NowcastNet and SimVP, reproduce the overall eastward motion but fail to capture the expansion of the western cell, resulting in fragmented and underdeveloped echoes. The parameter-free pySTEPS model underestimates the translation speed of the system and does not reproduce its area growth. In contrast, Nowcast3D more faithfully captures both the rapid eastward displacement and the structural expansion of the convective system.
The quantitative results support these visual assessments. Nowcast3D substantially outperforms all baseline models in neighborhood CSI, indicating more accurate localization of precipitation (Figure 3b). In addition, PSD and LPIPS evaluations show that its forecasts better reproduce the observed scale-dependent variability and exhibit higher perceptual similarity to the radar observations (Figure 3c,d). Because Nowcast3D produces ensembles, we also assess its probabilistic forecast skill. The Continuous Ranked Probability Score (CRPS) [matheson1976scoring, hersbach2000decomposition] shows that Nowcast3D yields better-calibrated forecasts than the ensemble-based NowcastNet, indicating more reliable uncertainty quantification (Figure 3e).
@@
Taking advantage of our model’s native 3D design, we further examine the vertical distribution of forecast CSI (Figure 3f). A layer-wise analysis shows that CSI is lower near the surface and at upper levels, but higher in the mid-troposphere. This pattern closely mirrors the vertical profile of mean reflectivity: echoes are strongest and most organized at mid-levels, leading to higher CSI, whereas near-surface and upper levels exhibit weaker and more intermittent echoes, naturally yielding lower CSI. The resulting vertical CSI profile is therefore closely tied to both the intrinsic storm structure and the radar sampling characteristics, rather than representing a simple ranking of forecast skill with height. In contrast, LPIPS shows much smaller inter-layer variation, indicating that the model maintains a consistent perceptual quality of forecasts across different heights. Layer-by-layer comparisons of the 3D forecasts are provided in Extended Data Figure 3. Additional precipitation case studies together with forecast comparisons are presented in Supplementary Note 6.3 (see Supplementary Figures 10–11 and Supplementary Videos 4–5).
@@
Figure 4: Aggregate quantitative evaluation of forecast skill. a, Demographics and institutional affiliations of the participating meteorologists, including gender and regional distribution. b, Evaluation of meteorologists’ model preferences. Compared with other competing baseline methods, meteorologists showed a stronger preference for Nowcast3D. c, Forecast skill metrics evaluated on the 0.04° resolution test set. The top panels show CSI as a function of lead time at 35 dBZ and 45 dBZ thresholds. The bottom-left panel shows PSD of forecasts compared with the ground truth (black dashed line), and the bottom-right panel shows the mean LPIPS scores, where lower values indicate better perceptual quality. d, The same set of metrics as in c, evaluated on the 0.01° resolution test set to assess the generalizability of different models at a finer spatial scale.
@@
To evaluate the practical utility of different models in forecasting extreme precipitation, we conducted an expert evaluation following the protocols established for DGMR [ravuri2021skilful] and NowcastNet [zhang2023skilful]. A group of senior forecasters across China were invited nationwide. Forecast animations from all models were displayed side by side, anonymized and presented in random order to minimize bias.
A total of 160 meteorologists from the China Meteorological Administration and 30 provincial or municipal observatories participated in the evalution. The demographic and regional distribution of participants is shown in Figure 4a. The experts cover nearly all regions of mainland China, except Tibet. Each meteorologist evaluated 23 test cases, randomly sampled from the precipitation events that occurred in China. Following the NowcastNet protocol [zhang2023skilful], we considered two complementary evaluation modes. In the post-hoc mode, forecasters ranked the model forecasts with access to the verifying radar observations. In the prior mode, they conducted ranking only using the preceding radar sequences, without access to future observations, to simulate real-time operational decision-making in the real world.
The aggregated preferences are summarized in Figure 4b. In the post-hoc evaluation, Nowcast3D was ranked best in 57% of cases, while, in the prior evaluation, it remained the top choice in 51% of cases. Overall, Nowcast3D received the highest preference among the participating meteorologists. These preferences are consistent with the model’s ability to represent the 3D evolution of reflectivity, enabled by its explicit treatment of physical fields and embedded dynamical constraints. In contrast, the 2D baseline models showed limitations in reproducing the full spatiotemporal structure of precipitation systems. A detailed description of the expert evaluation configuration is provided in Supplementary Note 4.
@@
To quantify the forecast skill, we employ three complementary metrics: neighbourhood CSI to assess location-dependent event detection, PSD to evaluate scale-dependent spatial structure, and LPIPS to measure perceptual and structural similarity. All metrics are computed over the full test corpus at two spatial scales: a provincial-scale domain at a standard 0.04° resolution and a more demanding urban domain at 0.01° resolution. This dual-scale design allows us to examine the model’s performance and robustness to finer spatial details.
At 0.04° resolution, Nowcast3D consistently outperforms all baseline models across all metrics (Figure 4c). For both moderate (35 dBZ) and intense (45 dBZ) reflectivity thresholds, it achieves the highest neighbourhood CSI throughout the 3-hour forecast horizon, indicating more accurate localization of heavy precipitation. The PSD analysis confirms that its forecasts most closely match the observed power spectrum, reflecting a more realistic representation of precipitation structures across spatial scales. This is corroborated by the LPIPS scores, where Nowcast3D attains the lowest values, indicating the highest perceptual and structural fidelity.
This advantage persists at the finer 0.01° resolution, which provides a stricter test of performance at the city scale (Figure 4d). Across all lead times and metrics, Nowcast3D maintains a clear margin over the competing baseline models, demonstrating that its architecture can be effectively adapted to finer spatial grids. Taken together, these results show that explicit modeling of 3D dynamics yields forecasts that are more accurate, structurally realistic and perceptually coherent than those produced by existing approaches. This benefit holds robustly across two different spatial scales.
Additional CSI curves over a wider range of intensity thresholds are provided in Supplementary Note 6.1 (see Supplementary Figure 5).
@@ Discussion
Precipitation nowcasting remains a central challenge in atmospheric science due to the strongly nonlinear, multiscale, and inherently non‑stationary nature of convective evolution, which contains a substantial stochastic component. Although recent data-driven advances have improved short-range guidance, skill for extreme precipitation still degrades rapidly along with lead time. Existing approaches primarily operate on 2D radar products that collapse the atmosphere into a single layer, discarding the height-dependent organization of convection and the layer-to-layer wind variations that shape storm motion and intensity through vertical motion and shear. Consequently, purely 2D methods often lack the physical fidelity needed for reliable, operationally useful nowcasts in hazardous weather regimes.
@@
Our work addresses these limitations by combining physics-informed dynamics with deep generative modeling on native 3D radar volumes. The core idea is establish a physics-informed operator network that learns to infer and extrapolate key physical drivers directly from sequences of volumetric reflectivity: a 3D wind field that governs advective transport, a local field that represents stochastic dispersion, and a residual tendency associated with unresolved microphysics. The forecast is advanced according to the learned operators, enforcing physical consistency and enabling physically plausible predictions at lead times. This physics-informed learning core is coupled to a conditional generative model that approximates the full distribution of future states, capturing uncertainty in both storm structure and localized intensity. Together, these components produce forecasts that are dynamically consistent and probabilistically reliable, delivering the uncertainty information required for risk-aware decision-making.
@@
The performance of this framework nonetheless depends on several external factors. First, its accuracy is constrained by the quality, coverage and scanning strategy of the input radar network. The beam attenuation, ground clutter and anomalous propagation also remain important sources of error. Second, inferring 3D flow from reflectivity alone is under-constrained, particularly in regimes with complex microphysics where phase changes can alter reflectivity without a corresponding change in the velocity field. These limitations point to clear directions for future work. Integrating additional observation types, such as Doppler radial velocities, geostationary satellite imagery and lightning measurements, could better constrain the inferred wind fields and improve the realism of hydrometeor evolution. Closer coupling with numerical weather prediction models, for example through data assimilation of the learned fields or by using NWP output as a physical prior, could further extend forecast skill beyond the traditional nowcasting range.
@@
In conclusion, by unifying physics-informed dynamics with deep generative modeling on 3D radar data, this study outlines a pathway toward developing new nowcasting systems. Such models go beyond simple pattern extrapolation to learn representations of the underlying physical processes, producing forecasts that are more accurate, reliable and physically consistent. Further enriching this framework with multimodal observations and tighter NWP coupling has potential to close the remaining gap towards robust, actionable forecasts of extreme precipitation events — a capability urgently needed in practice.
@@ Methods
Nowcast3D is a hybrid forecasting system that combines a physics-informed neural predictor with a diffusion-based generative model to produce probabilistic 3D radar reflectivity forecasts. As summarized in Extended Data Figure 1, the workflow has three stages. First, raw volumetric radar observations are passed through a data-processing module that derives a 3D spatial mask of observed and unobserved regions, applies mask-guided imputation only in the unobserved regions, and normalizes reflectivity to a fixed dynamic range. Second, a physics-informed network infers underlying physical fields from the recent radar history and advances them forward in time based on an explicit advection–diffusion–source equation, thereby producing a deterministic forecast. Third, a conditional diffusion model refines this physics-consistent forecast into an ensemble of stochastic realizations that quantify forecast uncertainty.
In this framework, the preprocessing stage defines the input reflectivity fields $R$ used throughout the paper. The physics-informed module implements a mapping, denoted by $\boldsymbol{\Theta}$, from the observed radar history to time-varying physical fields $X_{1:T}$ and a neural-numerical solver for the resulting evolution operator $\mathcal{F}$. The generative module implements a conditional diffusion model, denoted by $\boldsymbol{\Phi}$, that samples future trajectories conditioned on both past observations and the deterministic forecast. The following subsections describe the radar data and its preprocessing, the hybrid physics–generative formulation, the physics-informed forecasting module, the probabilistic refinement module, and the training procedure.
@@
Our forecasts are driven solely by 3D radar reflectivity from the national radar composite operated by the China Meteorological Administration. This product provides volumetric reflectivity on 24 constant-altitude levels between 0.5 and 16 km, with native horizontal resolutions down to 0.01°. These volumes resolve the 3D structure and evolution of precipitating systems over China and form the common input to all models evaluated in this study. Details of the dataset, including the storage format, are provided in Supplementary Note 1.1.
We consider three study domains that together span diverse convective regimes and spatial resolutions. The primary training domain is located in South China (110.00°–120.24°E, 23.96°–34.20°N), where we use all eligible radar volumes recorded in 2024. Cross-regional generalization at the native 0.04° resolution is assessed on a subset of precipitating scenes over North China (110.00°–120.24°E, 33.96°–44.20°N) observed in 2025. This subset is spatially and temporally disjoint from the training domain of South China. High-resolution coastal precipitation hazards are evaluated over the Maoming region in Guangdong Province, China (109.64°–112.20°E, 20.39°–22.95°N), where pretrained models at 0.04° resolution are fine-tuned on 0.01° Maoming sequences recorded in 2024. An independent subset of Maoming precipitation events in 2025 is reserved for testing. The radar composite coverage and the three study domains are shown in Extended Data Figure 5a. Note that Extended Data Figure 5b illustrates how terrain blocking and the shorter range of low-elevation scans lead to a pattern of sparse, near-surface circular observations.
Before inputting into the forecasting model, raw radar volumes pass through a preprocessing pipeline (see Extended Data Figure 1b). We first construct a mask that flags regions with missing measurements, including areas affected by terrain-induced beam blockage and range limitations. A non-learning, mask-guided imputation scheme is then applied to fill only the masked regions, using upper-level reflectivity as a predictor while leaving all observed echoes unchanged. Finally, reflectivity values are normalized to a fixed dynamic range. Extended Data Figure 5c,d compare raw and imputation reflectivity fields at 500 m and 1,000 m over the South China and Maoming regions, respectively. Further details on masking, imputation, and normalization are given in Supplementary Notes 1.2–1.4).
@@
Let $\{R_{t}\}_{t=-T_{0}}^{T}$ denote the preprocessed 3D radar reflectivity sequence, where each $R_{t}\in\mathbb{R}^{D\times H\times W}$. $D\times H\times W$ are the vertical and horizontal grid dimensions. The past observations are written as $R_{-T_{0}:0}\in\mathbb{R}^{(T_{0}+1)\times D\times H\times W}$. Given $R_{-T_{0}:0}$, our goal is to forecast the future sequence $R_{1:T}\in\mathbb{R}^{T\times D\times H\times W}$ using a unified two-stage framework that first produces a physics-based deterministic forecast and then refines it by a probabilistic ensemble model.
In the physics-informed learning stage, a neural network $\boldsymbol{\Theta}$ maps the historical radar sequence to a set of time-varying physical fields, $X_{1:T}=\boldsymbol{\Theta}(R_{-T_{0}:0})$, where $X_{1:T}\in\mathbb{R}^{T\times C\times D\times H\times W}$ and $C$ is the number of inferred quantities that govern the evolution of reflectivity. These fields act as coefficients in a time-dependent evolution equation, $\partial_{t}R=\mathcal{F}(R,X)$, where $\mathcal{F}$ represents the physical dynamics controlled by $X$. Starting from the initial condition $R(0)=R_{0}$ (the last observed volume), integrating this equation forward in time yields a deterministic forecast $R^{\mathrm{det}}_{1:T}$.
In the probabilistic forecasting stage, a conditional diffusion model generates an ensemble of forecasts around the physics-consistent trajectory produced by the deterministic core. The generative network $\boldsymbol{\Phi}$ is conditioned on the context $c=\{R_{-T_{0}:0},R^{\mathrm{det}}_{1:T}\}$, which combines the past observations with the deterministic prediction. Ensemble members are obtained by sampling the learned conditional distribution with different noise realisations $\epsilon^{(k)}$, $k=1,\dots,K$, thereby representing stochastic variability and forecast uncertainty.
The aforementioned steps can be written compactly as follows:
@@
Nowcast3D implements these formulations with a shared backbone that realizes $\boldsymbol{\Theta}$ and a conditional diffusion network that realizes $\boldsymbol{\Phi}$, as described in the following.
@@
We model the evolution of radar reflectivity in continuous time and space over a fixed forecast horizon. The deterministic forecast $R^{\mathrm{det}}_{1:T}$ is obtained by numerically integrating a partial differential equation. We specify the dynamics $\mathcal{F}(R,X)$ as an advection–diffusion–source equation, given by
@@
Here, the reflectivity field $R$ evolves under three processes: advective transport by a 3D velocity field $\boldsymbol{v}$; local anisotropic diffusion governed by a diffusivity tensor $\boldsymbol{\kappa}$; and a residual source term $s$ that summarizes unresolved processes, including microphysical growth and decay. The time-varying fields $(\boldsymbol{v},\boldsymbol{\kappa},s)$ over the forecast horizon constitute the physical fields $X_{1:T}$ and are inferred from the input radar sequence $R_{-T_{0}:0}$ by the physics module $\boldsymbol{\Theta}$.
Neural parameterization.
The time-varying physical fields $X_{1:T}=(\boldsymbol{v}_{1:T},\boldsymbol{\kappa}_{1:T},s_{1:T})$ required in Eq. (2) are produced by a shared encoder–decoder backbone followed by task-specific heads. We denote the backbone (encoder, U-Net branch, Transformer branch and fusion) by $\mathrm{Backbone}(\cdot)$. Its output is passed to three dedicated decoders that specialize in motion, diffusion and source components, expressed as follows: